limT(x-0){(a^x-1)/x then value of limit is

1
Expert's answer
2022-01-21T10:18:34-0500

Let "y=a^x-1, a>0," then "1+y=a^x," we have


"x=\dfrac{\ln(1+y)}{\ln a}""\lim\limits_{x\to 0}y=\lim\limits_{x\to 0}(a^x-1)=a^0-1=1-1=0"

Therefore, the given limit can be written as


"\lim\limits_{x\to 0}\dfrac{a^x-1}{x}=\lim\limits_{y\to 0}\dfrac{y}{\dfrac{\ln(1+y)}{\ln a}}"

"=\ln a\cdot\lim\limits_{y\to 0}\dfrac{1}{\ln(1+y)^{1/y}}=\dfrac{\ln a}{\ln (\lim\limits_{y\to 0}(1+y)^{1/y})}"

"=\dfrac{\ln a}{\ln (e)}=\ln a"

"\lim\limits_{x\to 0}\dfrac{a^x-1}{x}=\ln a"